OpenStudy (gabylovesyou):

What are the x-intercept(s) of the graph of y − x2 = 11x + 30? (−5, 0) and (6, 0) (5, 0) and (−6, 0) (−5, 0) and (−6, 0) (5, 0) and (6, 0)

OpenStudy (gabylovesyou):

@kausarsalley

OpenStudy (anonymous):

\[y − x^2 = 11x + 30\] \[y=x^2+11x+30\] \[x^2+11x+30=0\] and by some miracle this one factors as \[(x+5)(x+6)=0\] go from there

OpenStudy (anonymous):

rather i think it would be better this way y=(x+5)(x+6)

OpenStudy (gabylovesyou):

C...

OpenStudy (gabylovesyou):

right ? @satellite73

OpenStudy (anonymous):

yes.....that is it

OpenStudy (gabylovesyou):

thanks :p

OpenStudy (anonymous):

you are welcome....